Slope=-2, passes through (4,1)
Found 2 solutions by John10, shree840:
Answer by John10(297)

Slope m = -2, passes through (4,1)
Use the slope -point formula to write the equation:
y-yo = m(x-xo)
y - 1 = -2(x - 4)
y -1 = -2x + 8
y = -2x + 9
Have a great day!John10

Answer by shree840(260)

slope=-2 passing through (4,1)
y=mx+b
Y=-2x+b Passing through 4,1 substitute these values to get b
1=-2*4+b=-8+b or b=9
equation is y=-2x+9 ANS